A particle executes simple harmonic motion with an amplitude of 10 cm. At what distance from the mean position are the kinetic and potential energies equal ?


 

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello
for any particle
x=10sin(2pi f t)
v=10*2pi f cos(2pi f t)
1/2mv^2=mgh
=>v^2=2gx
Please put in the value to get phase of the wave
Hint (use cos^2=1-sin^2 to solve the equation)
thats the answer
Thanks & Regards
Arun Kumar
Btech IIT delhi
Askiitians Faculty